Twatalika acquired a new property (land and buildings) in Kaoma, on 1 January 2018 for K40 million (including K15 million in respect of the land). The asset was revalued on 31 December 2019 to K43 million (including K16.6 million in respect of the land). The building’s element will be depreciated over a 50-year useful life with a residual value of nil. The useful life and residual value did not subsequently need revision. On 31 December 2020, the property was revalued downwards to a fair value of K35 million as a result of the recession (including K14 million in respect of the land). The company makes the annual transfers from revaluation surplus to retained earnings in respect of excess depreciation, as allowed by IAS 16 Property, Plant, and Equipment. Calculate the amounts that must be reported in the profit or loss and other comprehensive income in respect of the building for the years ended 31 December 2019 and 31 December 2020.

1
Expert's answer
2022-03-13T18:59:23-0400

2019 - depreciation = 25m/50 = 0.5m

2020 - depreciation = 36.4/48 = 0.75m

revaluation gain in 2019

land = 16.6- 15 = 1.6m

building = 36.4 - (25-(0.5+0.5)) = 36.4-24 = 12.4m

Total revaluation gain = 1.6+12.4 = 14m

Revaluation 2020

land = 14 -16.6 = -2.6

Building = 21 - (36.4- 0.75) = -14.65

Total revaluation Loss = 17.25

Charges in Profit and Loss account and comprehensive a/c

2019

Depreciation 0.5m

comprehensive income 14m


Charges in Profit and Loss account and comprehensive a/c

2020

Depreciation 0.75m

comprehensive income (loss) 17.25m